### v2

# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

Abstract
It has been known that stochastic gravitational wave backgrounds (SGWBs) have anisotropies generated by squeezed-type tensor non-Gaussianities originating from scalar-tensor-tensor (STT) and tensor-tensor-tensor cubic interactions. While the squeezed tensor non-Gaussianities in the standard slow-roll inflation with the Bunch-Davies vacuum state are suppressed due to the so-called consistency relation, those in extended models with the violation of the consistency relation can be enhanced. Among such extended models, we consider the inflation model with the non-Bunch-Davies state that has been known to enhance the squeezed tensor non-Gaussianities. We explicitly formulate the primordial STT bispectrum induced during inflation in the context of Horndeski theory with the non-Bunch-Davies state and show that the induced SGWB anisotropies can be enhanced. We then discuss the detectability of those anisotropies in future gravitational wave experiments.
1 Introduction
Recently, several international collaborations for the Pulsar Timing Array (PTA) gravitational wave (GW) experiments such as North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [1, 2] have reported evidence of nHz stochastic gravitational wave background (SGWB) [3, 4, 5, 6]. Furthermore, future ground- and space-based missions at higher frequency bands, e.g., Cosmic Explorer (CE) [7], Einstein Telescope (ET) [8], Laser Interferometer Space Antenna (LISA) [9, 10], Taiji [11], the Taiji-LISA network [12, 13], Decihertz Gravitational Wave Observatory (DECIGO) [14, 15], the Big Bang Observer (BBO) [16], etc., are expected to report other evidence for the SGWB at different frequency ranges. While the most plausible source of the SGWB is the superposition of the GWs from astrophysical compact objects such as black holes (see, e.g., Ref. [17]), various cosmological origins of the SGWB, in particular, from the early universe with new physics have also been widely discussed (see, e.g., Ref. [18]). It is, therefore, worth discussing the possibilities of specifying the source of SGWB and testing new physics in the early universe with the PTA experiments and the ground- and space-based missions.
As one of the things that can help identify the sources, the anisotropies of SGWB have been recently investigated [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. In addition to the anisotropies associated with the metric fluctuations in the late time universe (see, e.g., Refs. [19, 20, 24, 26]) known as the so-called Sachs-Wolfe (SW) and integral SW effects for the CMB anisotropies, [^1] See also Refs. [25, 31, 33, 34, 35, 36] where the anisotropies of secondary gravitational waves sourced by the late-time perturbation have been studied. the SGWB anisotropies can be sourced by the primordial scalar-tensor-tensor (STT) and tensor-tensor-tensor (TTT) cubic interactions in the squeezed configuration [21, 22, 23, 27, 28, 29, 30, 32]. In the present paper, we discuss the SGWB anisotropies induced by the primordial STT correlations generated during inflation. In the standard single scalar inflation, the consistency relation prohibits any enhancements of the primordial non-Gaussianities in the squeezed configuration [37]. On the other hand, in extended models with its violation, those non-Gaussianities can be enhanced, and then the SGWB anisotropies in such non-standard inflationary models have been studied in Refs. [21, 22, 23, 27, 28, 29, 30, 32]. Here, we consider inflation with non-Bunch-Davies states among the extended models with the consistency relation being violated.
Inflation with the non-Bunch-Davies states has been shown to potentially enhance the primordial non-Gaussianities as a consequence of the violation of the consistency relation [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. In particular, the bispectra of the scalar and tensor perturbations are enhanced in proportion to powers of $k_{S}/k_{L}$ where $k_{S}$ and $k_{L}$ respectively denote the comoving wavenumber of the short- and long-wavelength modes in the squeezed limit, that is, $k_{L}/k_{S}\to 0$, and then it is naively expected that such a $k_{S}/k_{L}$ enhancement could efficiently amplify the SGWB anisotropies. However, since this characteristic enhancement of the primordial non-Gaussianities due to the non-Bunch-Davies states is basically induced by the cubic interactions on subhorizon scales, the short-wavelength modes may suffer from a so-called Trans-Planckian problem [79] in the squeezed limit $k_{S}\gg k_{L}$ when $k_{S}$ and $k_{L}$ are respectively taken to be the target frequency of observed SGWB and the scale of interest for their anisotropies on large scales. To obtain proper predictions for the SGWB anisotropies from the inflation with non-Bunch-Davies states, we first derive a condition to avoid the Trans-Planckian problem and then obtain the primordial STT bispectrum from the inflation with non-Bunch-Davies states under that condition. Also, towards testing the SGWB anisotropies from inflation, it is convenient to study those in a general framework of inflation, and hence we work with the Horndeski theory which is the most general scalar-tensor theory whose equations of motion are second order [80, 81, 82] (see also Ref. [83] for a review). As a result, we will show that the resultant non-Gaussianities and SGWB anisotropies from the general single-field inflation can be enhanced in a different way than the $k_{S}/k_{L}$ enhancement. We also discuss whether the SGWB anisotropies can be detected with actual gravitational wave experiments. The DECIGO/BBO experiments are designed to test inflation models predicting an almost scale-invariant tensor spectrum with the tensor-to-scalar ratio being $\mathcal{O}(10^{-2})$, while the recent NANOGrav data can be explained from inflation with a certain blue-tilted tensor spectrum. The power spectrum of the non-Bunch-Davies tensor modes generally has scale dependence via a Bogoliubov coefficient, and hence it is possible for inflation with the non-Bunch-Davies state to explain the recent results of nHz SGWB [84]. In light of these GW observations, we consider two configurations of the Bogoliubov coefficient. One is a scale-independent coefficient. This configuration leads to an almost scale-invariant tensor spectrum, the SGWB originating from which can potentially be tested by DECIGO/BBO. The other is a coefficient peaked at a certain scale $k=k_{*}$. The resultant tensor power spectrum has a peak at $k=k_{*}$ as well, and hence, properly choosing the value of $k_{*}$, the resultant GW can explain the recent PTA data. We would also like to investigate the cross-correlation between the cosmic microwave background (CMB) and the SGWB anisotropies [27].
This paper is organized as follows. In the following section, we briefly review our setup and primordial power spectra of scalar and tensor perturbations with the non-Bunch-Davies states. In Sec. 3, we first explain the Trans-Planckian problem for the short-wavelength GWs and clarify the condition to avoid it. Then, we compute the primordial STT bispectrum with the non-Bunch-Davies states under that condition. In Sec. 2.2, we introduce two phenomenological models for the Bogoliubov coefficient such that the tensor power spectrum is nearly scale-invariant or has a peak at a certain scale. In Sec. 4, after reviewing the impacts of the primordial non-Gaussianities on the SGWB anisotropies, we evaluate the auto-correlation function of those and the cross-correlation one between those and CMB anisotropies. Then, we discuss the detectability of those correlation functions. The summary of this paper is drawn in Sec. 5.
2 Primordial power spectra with non-Bunch-Davies states
To explore the impact of primordial non-Gaussianities induced by the non-Bunch-Davies states on the SGWB, it is useful to formulate those in a general framework of single-field inflation. We work with the Horndeski theory [80, 81, 82], which is the most general scalar-tensor theory whose equations of motion are second order. The action takes the following form [82],
$S$ $=\int{\rm d}^{4}x\sqrt{-g}\biggl{[}G_{2}(\phi,X)-G_{3}(\phi,X)\Box\phi+G_{4}(\phi,X)R+G_{4X}[(\Box\phi)^{2}-(\nabla_{\mu}\nabla_{\nu}\phi)^{2}]$
$\quad+G_{5}(\phi,X)G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\phi-\frac{1}{6}G_{5X}[(\Box\phi)^{3}-3\Box\phi(\nabla_{\mu}\nabla_{\nu}\phi)^{2}+2(\nabla_{\mu}\nabla_{\nu}\phi)^{3}]\biggr{]},$ (2.1)
with $G_{i}(\phi,X)$ being arbitrary functions of the scalar field $\phi$ and its canonical kinetic term $X:=-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi/2$. Imposing appropriate conditions on the arbitrary functions, the scalar field drives the slow-roll inflation as an inflaton [82].
The perturbed metric around a spatially flat Friedmann-Lemaître-Robertson-Walker spacetime under the unitary gauge in which the spatial fluctuation of the scalar field vanishes, that is, $\phi=\phi(t)$, is of the form
${\rm d}s^{2}=-N^{2}{\rm d}t^{2}+g_{ij}({\rm d}x^{i}+N^{i}{\rm d}t)({\rm d}x^{j}+N^{j}{\rm d}t),$ (2.2)
where
$N$ $=1+\alpha,\ N_{i}=\partial_{i}\beta,$ (2.3)
$g_{ij}$ $=a^{2}e^{2\zeta}\biggl{(}\delta_{ij}+h_{ij}+\frac{1}{2}h_{ik}h_{kj}+\cdots\biggr{)},$ (2.4)
with $a(t)$ being a scale factor. Here, $\alpha,~{}\beta,$ and $\zeta$ are scalar perturbations, and $h_{ij}$ means transverse-traceless tensor perturbations. We hereafter use a dot as a derivative with respect to $t$, and the Hubble parameter is defined by $H:=\dot{a}/a$.
2.1 Quadratic action and primordial power spectra with non-Bunch-Davies states
Expanding the action Eq. (2.1) up to the quadratic order in the perturbations and substituting the solutions of $\alpha$ and $\beta$ derived from the constraint equations into the perturbed action, one obtains [82]
$S^{(2)}_{\zeta}$ $=\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}\mathcal{G}_{S}\dot{\zeta}^{2}-\frac{\mathcal{F}_{S}}{a^{2}}(\partial_{i}\zeta)^{2}\biggr{]},$ (2.5)
$S^{(2)}_{h}$ $=\frac{1}{8}\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}\mathcal{G}_{T}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr{]},$ (2.6)
where $\mathcal{G}_{S},~{}\mathcal{F}_{S},~{}\mathcal{G}_{T}$, and $\mathcal{F}_{T}$ are given in terms of $\phi$, $X$, the arbitrary functions $G_{i}(\phi,X)$ and their derivatives with respect to $\phi$ and $X$, and the explicit expressions of these functions are given in Appendix A. Throughout this paper, we assume a slow-roll inflation with $\mathcal{G}_{S},~{}\mathcal{F}_{S},~{}\mathcal{G}_{T},~{}\mathcal{F}_{T}\simeq{\rm const.}$ in a quasi-de Sitter spacetime ($H\simeq{\rm const.}$). Specifically, we introduce the slow-roll parameters as
$\epsilon:=-\frac{\dot{H}}{H^{2}},\ g_{S}:=\frac{\dot{\mathcal{G}}_{S}}{H\mathcal{G}_{S}},\ f_{S}:=\frac{\dot{\mathcal{F}}_{S}}{H\mathcal{F}_{S}},\ g_{T}:=\frac{\dot{\mathcal{G}}_{T}}{H\mathcal{G}_{T}},\ f_{T}:=\frac{\dot{\mathcal{F}}_{T}}{H\mathcal{F}_{T}},$ (2.7)
and assume $\epsilon,~{}g_{S},~{}f_{S},~{}g_{T},~{}f_{T}\ll 1$ as slow-roll approximations.
We then quantize the scalar and tensor perturbations. The Fourier transformations of those are defined by
$\zeta(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{\zeta}(t,{\bf k})e^{i\bf k\cdot\bf x},$ (2.8)
$h_{ij}(t,{\bf x})$ $=\int\frac{{\rm d}^{3}k}{(2\pi)^{3}}\tilde{h}_{ij}(t,{\bf k})e^{i\bf k\cdot\bf x}.$ (2.9)
In Fourier space, both perturbations are quantized as
$\tilde{\zeta}\to~{}\hat{\zeta}(t,{\bf k})$ $=u_{k}(t)\hat{a}_{\bf k}+u_{k}^{*}(t)\hat{a}^{\dagger}_{-{\bf k}},$ (2.10)
$\tilde{h}_{ij}\to~{}\hat{h}_{ij}(t,{\bf k})$ $=\sum_{s}\biggl{[}v_{k}(t)e^{(s)}_{ij}({\bf k})\hat{a}^{(s)}_{\bf k}+v^{*}_{k}(t)e^{(s)*}_{ij}(-{\bf k})\hat{a}^{(s)\dagger}_{-{\bf k}}\biggr{]},$ (2.11)
where $e^{(s)}_{ij}(\bf k)$ denotes the polarization tensor with the helicity states $s=\pm 2$, satisfying the transverse-traceless conditions, i.e., $k_{i}e^{(s)}_{ij}({\bf k})=\delta_{ij}e^{(s)}_{ij}({\bf k})=0$. The creation and annihilation operators satisfy the standard commutation relations,
$[\hat{a}_{\bf k},\hat{a}^{\dagger}_{{\bf k}^{\prime}}]$ $=(2\pi)^{3}\delta({\bf k}-{\bf k}^{\prime}),$ (2.12)
$[\hat{a}^{(s)}_{\bf k},\hat{a}^{(s^{\prime})\dagger}_{{\bf k}^{\prime}}]$ $=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}-{\bf k}^{\prime}),$ (2.13)
${\rm others}$ $=0.$ (2.14)
The equations of motion for the perturbations in real space are of the form,
$E^{s}$ $:=\partial_{t}(a^{3}\mathcal{G}_{S}\dot{\zeta})-a\mathcal{F}_{S}\partial^{2}\zeta=0,$ (2.15)
$E^{h}_{ij}$ $:=\partial_{t}(a^{3}\mathcal{G}_{T}\dot{h}_{ij})-a\mathcal{F}_{T}\partial^{2}h_{ij}=0.$ (2.16)
On the quasi-de Sitter background and under the slow-roll approximations, at the leading order, general solutions of the equations of motion for the mode functions can be obtained as
$u_{k}$ $=\alpha_{k}u_{k,({\rm BD})}+\beta_{k}u^{*}_{k,({\rm BD})},$ (2.17)
$v^{(s)}_{k}$ $=\alpha^{(s)}_{k}v_{k,({\rm BD})}+\beta^{(s)}_{k}v^{*}_{k,({\rm BD})},$ (2.18)
where $\alpha_{k},~{}\beta_{k},~{}\alpha^{(s)}_{k}$, and $\beta^{(s)}_{k}$ are Bogoliubov coefficients, and $u_{k,({\rm BD})}$ and $v_{k,({\rm BD})}$ denote the mode functions with the Bunch-Davies state which are respectively given as [82]
$u_{k,({\rm BD})}$ $=\frac{iH}{2\sqrt{\mathcal{F}_{S}c_{s}k^{3}}}(1+ic_{s}k\eta)e^{-ic_{s}k\eta},$ (2.19)
$v_{k,({\rm BD})}$ $=\frac{i\sqrt{2}H}{\sqrt{\mathcal{F}_{T}c_{h}k^{3}}}(1+ic_{h}k\eta)e^{-ic_{h}k\eta},$ (2.20)
where we introduced the conformal time $\eta:=\int{\rm d}t/a$ and the propagation speed of each mode $c_{s}^{2}=\mathcal{F}_{S}/\mathcal{G}_{S},~{}c_{h}^{2}=\mathcal{F}_{T}/\mathcal{G}_{T}$. The Bogoliubov coefficients satisfy the following normalization conditions:
$|\alpha_{k}|^{2}-|\beta_{k}|^{2}$ $=1,$ (2.21)
$|\alpha^{(s)}_{k}|^{2}-|\beta^{(s)}_{k}|^{2}$ $=1,$ (2.22)
which are derived from the Wronskian conditions. The dimensionless primordial power spectra of the scalar and tensor perturbations are respectively defined by
$\langle\hat{\zeta}({\bf k})\hat{\zeta}({\bf k}^{\prime})\rangle=(2\pi)^{3}\delta({\bf k}+{\bf k}^{\prime})\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{\zeta},$ (2.23)
$\langle\hat{\xi}^{(s)}({\bf k})\hat{\xi}^{(s^{\prime})}({\bf k}^{\prime})\rangle=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}+{\bf k}^{\prime})\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h},$ (2.24)
where $\hat{\xi}^{(s)}({\bf k}):=\hat{h}_{ij}({\bf k})e^{(s)*}_{ij}({\bf k})$. Then, the power spectra on superhorizon scales ($-c_{s}k\eta\ll 1$ and $-c_{h}k\eta\ll 1$), can be obtained from the mode functions as
$\mathcal{P}_{\zeta}$ $=\frac{1}{8\pi^{2}}\frac{H^{2}}{\mathcal{F}_{S}c_{s}}|\alpha_{k}-\beta_{k}|^{2},$ (2.25)
$\mathcal{P}^{(s)}_{h}$ $=\frac{1}{\pi^{2}}\frac{H^{2}}{\mathcal{F}_{T}c_{h}}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (2.26)
Then, up to the leading order in the slow-roll parameters, the spectral indices of the scalar and tensor power spectra are derived as
$n_{s}-1$ $:=\frac{{\rm d}\ln\mathcal{P}_{\zeta}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{S}}{2}-\frac{3}{2}f_{S}+\frac{{\rm d}}{{\rm d}\ln k}\ln|\alpha_{k}-\beta_{k}|^{2},$ (2.27)
$n_{t}$ $:=\frac{{\rm d}\ln\mathcal{P}_{h}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{T}}{2}-\frac{3}{2}f_{T}+\frac{{\rm d}}{{\rm d}\ln k}\ln\sum_{s}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (2.28)
where $\mathcal{P}_{h}=\sum_{s}\mathcal{P}^{(s)}_{h}$. By taking the limit, $\alpha_{k}=1,~{}\beta_{k}=0,~{}\alpha^{(s)}_{k}=1$, and $\beta^{(s)}_{k}=0$, Eqs. (2.25)–(2.28) reproduce the results in the slow-roll inflation with the Bunch-Davies state [82]. In the main text, we hereafter set $\alpha_{k}=1$ and $\beta_{k}=0$ (i.e., assume the Bunch-Davies scalar modes) for simplicity, and we solely study the impact of the non-Bunch-Davies tensor modes on the SGWB and its anisotropies. (See Ref. [85] for the impacts of primordial gravitational waves induced by non-Bunch-Davies scalar modes on the SGWB.)
2.2 On the scale-dependence of the tensor power spectrum with non-Bunch-Davies state
The last term in Eq. (2.28) is peculiar to the excited tensor modes, and $k$ dependence of the Bogoliubov coefficients can yield a variety of non-standard predictions, e.g., a highly blue-tilted spectrum [86, 84]. Such a blue-tilted spectrum can potentially be tested by PTA experiments such as, e.g., NANOGrav, as has been studied in Ref. [84] where a power-law $k$ dependence of the power spectrum was assumed. On the other hand, the nearly-scale-invariant spectrum also has a chance to be tested by planned space interferometers such as DECIGO/BBO. Thus, here, we introduce phenomenological models for the Bogoliubov coefficients to realize sufficient amplitude of the SGWB to detect in both frequency bands. (See, e.g., Refs. [87, 88, 89, 90, 91] where impacts of non-Bunch-Davies states on scalar/tensor power spectra have been studied with different Bogoliubov coefficients.)
2.2.1 Peaked spectrum
To explore the possible enhancement at a certain scale, we consider an example in which the primordial tensor power spectrum has a peak at a scale $k=k_{*}$. Let us first impose the following ansatz on the Bogoliubov coefficients:
$\alpha^{(s)}_{k}$ $=1+i{\rm Im}[\alpha^{(s)}_{k}],$ (2.29)
$\beta^{(s)}_{k}$ $={\rm Re}[\beta^{(s)}_{k}].$ (2.30)
From the normalization condition Eq. (2.22), one can derive
$({\rm Im}[\alpha^{(s)}_{k}])^{2}=({\rm Re}[\beta^{(s)}_{k}])^{2}.$ (2.31)
We hereafter assume ${\rm Im}[\alpha^{(s)}_{k}]={\rm Re}[\beta^{(s)}_{k}]$ for simplicity. Then, to realize the peak, we further consider
${\rm Re}[\beta^{(s)}_{k}]=A\times\exp[-\frac{(\ln{k/k_{*}})^{2}}{\Delta^{2}}\biggr{]},$ (2.32)
where $A$ and $\Delta$ are dimensionless parameters that control how large and how sharp the peak is. For $A=-3000$ and $\Delta=0.45$, we can realize a blue-tilted tensor spectrum with $n_{t}\simeq 2$ for $k\lesssim k_{*}$, and an $\mathcal{O}(10^{7})$ enhancement in the power spectrum at a reference scale $k=k_{*}$. Ref. [84] mentioned that the nHz SGWB detected by recent PTA experiments [92, 93, 94, 4, 95, 96, 97, 98, 99, 100] can be explained by the single-field inflation with the non-Bunch-Davies vacuum state with choosing $k_{*}\sim 10^{8}{\rm Mpc}^{-1}$.
2.2.2 Nearly-scale-invariant spectrum
For a nearly-scale-invariant case, we employ the following ansatz: [^2] Here, we assumed a pure imaginary $\beta^{(s)}_{k}$. This is because the scale-invariant case with a pure real $\beta^{(s)}_{k}$ yields a leading-order contribution of $\mathcal{O}(B^{2})$ in Eqs. (4.48) and (4.49). As will be explained later, one has $B\ll 1$ due to the backreaction constraint. Therefore, we need a pure imaginary $\beta^{(s)}_{k}$ that leads to the leading-order contribution of $\mathcal{O}(B)$ in Eqs. (4.48) and (4.49).
$\alpha^{(s)}_{k}$ $=1+iB,$ (2.33)
$\beta^{(s)}_{k}$ $=iB,$ (2.34)
where $B$ is real and constant. In this case, the last term in Eq. (2.28) is zero, and the spectral index of the tensor power spectrum is given by the slow-roll parameters.
3 Non-Bunch-Davies bispectrum
As we will see later in section 4, the SGWB anisotropies on large scales are sensitive to the primordial non-Gaussianity which gives the coupling between short- and long-wavelength modes (see, e.g., Ref. [29, 32]). At the leading order, such a coupling can be characterized by the primordial bispectrum in the squeezed limit. Although the tensor auto-bispectrum in the squeezed limit can be a source of the anisotropies of primordial GWB, due to the tight observational constraint on the primordial tensor modes on CMB scales the contribution from the tensor auto-bispectrum might be smaller than that from the STT cross-bispectrum. Then, here, we focus on the primordial STT bispectrum in the squeezed limit, that is, the coupling between long-wavelength scalar modes and short-wavelength tensor modes.
3.1 Trans-Planckian problem
Before formulating the STT bispectrum, we mention the Trans-Planckian issue [79] in the inflation models with the non-Bunch Davies state. In general, the physical momenta of short-wavelength modes, e.g., GWs targeted by PTA such as NANOGrav or space interferometer such as DECIGO and BBO, can exceed the Planck scale (i.e., $k_{\rm phys}:=k/a\geq M_{\rm pl}$) when both short- and long-wavelength modes are on subhorizon scales at the same time. The condition that the long-wavelength modes are on subhorizon scales at $\eta=\eta_{0}$ is expressed as
$1<-k_{L}\eta_{0}\,\left(\,\ll-k_{S}\eta_{0}\,\right),$ (3.1)
where $k_{S}$ and $k_{L}$ denote the comoving wavenumber of the short- and long-wavelength modes, respectively. In the quasi-de Sitter background (i.e., $a\simeq-(H\eta)^{-1}$), Eq. (3.1) turns into
$\frac{k_{{\rm phys},L}}{M_{\rm pl}}>\frac{H_{\rm inf}}{M_{\rm pl}},$ (3.2)
and for the short-wavelength modes
$\frac{k_{{\rm phys},S}}{M_{\rm pl}}>\frac{H_{\rm inf}}{M_{\rm pl}}\times\frac{k_{S}}{k_{L}}.$ (3.3)
Assuming $\mathcal{O}(\mathcal{G}_{T},\mathcal{F}_{T})=M_{\rm Pl}^{2}$ and $|\alpha^{(s)}_{k_{\rm CMB}}-\beta^{(s)}_{k_{\rm CMB}}|^{2}\sim 1$, Eq. (2.26) gives $\mathcal{P}_{h}^{1/2}|_{k_{\rm CMB}}\sim H_{\rm inf}/M_{\rm pl}$, and then we have
$10^{-6}\left(\frac{r|_{k_{\rm CMB}}}{10^{-3}}\right)^{1/2}\left(\frac{\mathcal{P}_{\zeta}|_{k_{\rm CMB}}}{2\times 10^{-9}}\right)^{1/2}\left(\frac{k_{S}}{k_{L}}\right)<\frac{k_{{\rm phys},S}}{M_{\rm pl}},$ (3.4)
where $r:=\mathcal{P}_{h}/\mathcal{P}_{\zeta}$ is the tensor-to-scalar ratio. It is expected that future CMB B-mode polarization experiments can detect the inflationary SGWB with $r|_{k_{\rm CMB}}\sim 10^{-3}$, and then the above inequality means that the detection of the inflationary SGWB would suffer from the Trans-Planckian problem of the short-wavelength modes at $\eta=\eta_{0}$ for $k_{S}>10^{4}~{}{\rm Mpc}^{-1}\times(k_{L}/(10^{-2}~{}{\rm Mpc}^{-1}))$. Primordial non-Gaussianities from inflation with the Bunch-Davies state are insensitive to subhorizon modes since those are mainly generated after their horizon-cross scales, while those with the non-Bunch-Davies state are not. This is because the excited modes on subhorizon scales can generate the non-Gaussianities [39] and then the above Trans-Planckian issue becomes serious.
In principle, to avoid this problem one can consider a non-standard model of inflation with $\mathcal{P}_{\zeta}|_{k_{\rm CMB}}\sim 10^{-9}$ and $r\ll 10^{-3}$ by choosing model parameters, but we do not consider such models. Instead, we relax the starting point of the above argument Eq. (3.1) in such a way that the tensor modes got excited from the Bunch-Davies state at $\eta=\eta_{0}$ when the long-wavelength modes of $k_{L}=10^{-2}{\rm Mpc}^{-1}$ were already on superhorizon scales, while the short-wavelength modes were still on subhorizon scales:
$-k_{S}\eta_{0}>1\gg-k_{L}\eta_{0},$ (3.5)
and hence we obtain
$\frac{k_{{\rm phys},S}}{M_{\rm pl}}>r^{1/2}\mathcal{P}^{1/2}_{\zeta}|_{k_{\rm CMB}}\gg\frac{k_{{\rm phys},L}}{M_{\rm pl}}.$ (3.6)
By choosing $\eta_{0}$ appropriately, the Trans-Planckian problem can be avoided even for $k_{S}\gg 10\,{\rm Mpc}^{-1}$, e.g., the modes corresponding to the ones targeted by NANOGrav ($10^{6}\,\rm Mpc^{-1}$) and DECIGO/BBO ($10^{13}\,\rm Mpc^{-1}$). It has been extensively discussed that the non-Bunch-Davies tensor modes can enhance primordial bispectra compared to those in the case of the Bunch-Davies state. In particular, the enhancements of the so-called squeezed bispectra are in proportion to powers of $k_{S}/k_{L}$ (see, e.g., Ref. [51]). However, we will show that this enhancement appears in the case with Eq. (3.1), while it does not with Eq. (3.5).
3.2 Scalar-tensor-tensor bispectrum
The scalar-tensor-tensor bispectrum is defined by
$\langle\hat{\zeta}({\bf k}_{1})\hat{\xi}^{(s_{2})}({\bf k}_{2})\hat{\xi}^{(s_{3})}({\bf k}_{3})\rangle=(2\pi)^{3}\delta({\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3})\mathcal{B}_{shh}^{s_{2}s_{3}}.$ (3.7)
By expanding the action Eq. (2.1) up to the cubic order in the perturbations, the interaction Hamiltonian involving one scalar and two tensors is obtained as [101]
$H_{\rm int}=-\int{\rm d}^{3}x\mathcal{L}_{shh},$ (3.8)
with
$\mathcal{L}_{shh}$ $=a^{3}\biggl{[}b_{1}\zeta\dot{h}_{ij}^{2}+\frac{b_{2}}{a^{2}}\zeta(\partial_{k}h_{ij})^{2}+b_{3}(\partial_{k}\psi)\dot{h}_{ij}\partial_{k}h_{ij}$
$\qquad\qquad+b_{4}\dot{\zeta}\dot{h}_{ij}^{2}+\frac{b_{5}}{a^{2}}\partial^{2}\zeta\dot{h}_{ij}^{2}+b_{6}\psi_{,ij}\dot{h}_{ik}\dot{h}_{jk}+\frac{b_{7}}{a^{2}}\zeta_{,ij}\dot{h}_{ik}\dot{h}_{jk}\biggr{]}.$ (3.9)
As is summarized in Appendix A, the above cubic Lagrangian includes other terms proportional to the equations of motion for the linear perturbations which, in principle, can contribute to the bispectrum. However, those are always subleading under the slow-roll approximations, and hence we do not consider those in the main text.
For later convenience, we parametrize the primordial bispectrum as
$\mathcal{B}_{shh}^{s_{2}s_{3}}$ $={\rm Re}\left[\sum^{7}_{i=1}\left(\alpha^{(s_{2})}_{k_{2}}-\beta^{(s_{2})}_{k_{2}}\right)\left(\alpha^{(s_{3})}_{k_{3}}-\beta^{(s_{3})}_{k_{3}}\right)\frac{2}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\frac{H^{6}}{\mathcal{F}_{S}\mathcal{F}_{T}^{2}c_{s}c_{h}^{2}}b_{i}\mathcal{F}_{b_{i}}\mathcal{V}_{b_{i}}+(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\right].$ (3.10)
The functions $\mathcal{V}_{b_{i}}$ originate from products of the polarization tensors, the explicit expressions of which are summarized in Appendix B. To compute the primordial bispectrum, we perform the time integral of the in-in formalism from $\eta=\eta_{0}$ to the end of inflation, $\eta=0$. In general, the time integral can be schematically expressed as
$S(\tilde{K}):=\int_{\eta_{0}}^{0}{\rm d}\eta(-\eta)^{n}e^{-i\tilde{K}\eta},$ (3.11)
where $\tilde{K}$ takes $K^{\prime},K_{1}^{\prime},K_{2}^{\prime}$, or $K_{3}^{\prime}$ and they are respectively defined as
$K^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (3.12)
$K_{1}^{\prime}$ $:=-c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (3.13)
$K_{2}^{\prime}$ $:=c_{s}k_{1}+c_{h}(-k_{2}+k_{3}),$ (3.14)
$K_{3}^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}-k_{3}).$ (3.15)
Under Eq. (3.5) with $k_{1}=k_{L}$ and $k_{2}\simeq k_{3}\simeq k_{S}$, one can obtain
$S(\tilde{K})=\begin{cases}\displaystyle\mathcal{O}(\tilde{K}^{-(n+1)})&\ \ (\tilde{K}=K^{\prime}\ {\rm or}\ K_{1}^{\prime}),\\ \displaystyle\mathcal{O}(\eta_{0}^{(n+1)})&\ \ (\tilde{K}=K_{2}^{\prime}\ {\rm or}\ K_{3}^{\prime}).\end{cases}$ (3.16)
It is worth mentioning the difference of $S(\tilde{K})$ between Eq. (3.1) and Eq. (3.5). In the former, one has $S(\tilde{K})=\mathcal{O}(\tilde{K}^{-(n+1)})$ for all of the four cases under the squeezed limit, which has resulted in the enhancements of the squeezed bispectrum in proportion to powers of $k_{S}/k_{L}$. In the latter, one has $S(\tilde{K})=\mathcal{O}(\eta_{0}^{(n+1)})$, which will result in the enhancements in proportion to powers of $-k_{S}\eta_{0}$ as will be shown below. Since we have $k_{S}/k_{L}\gg-k_{S}\eta_{0}$, the enhancements will be reduced compared to the results in the literature (e.g., Ref. [51]) where Eq. (3.1) was assumed.
We summarize the details of computations in Appendix B, and we here show only the final results. The primordial STT bispectrum obtained under Eq. (3.5) is as follows,
$\mathcal{F}_{b_{1}}$ $=\frac{c_{h}^{4}k_{2}^{2}k_{3}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{1},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.17)
$\mathcal{F}_{b_{2}}$ $=\frac{1}{H^{2}}\biggl{[}\mathcal{I}_{b_{2},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.18)
$\mathcal{F}_{b_{3}}$ $=\frac{c_{h}^{2}c_{s}^{2}k_{1}^{2}k_{2}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{3},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.19)
$\mathcal{F}_{b_{4}}$ $=\frac{c_{h}^{4}c_{s}^{2}k_{1}^{2}k_{2}^{2}k_{3}^{2}}{H}\biggl{[}\mathcal{I}_{b_{4},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.20)
$\mathcal{F}_{b_{5}}$ $=c_{h}^{4}k_{1}^{2}k_{2}^{2}k_{3}^{2}\biggl{[}\mathcal{I}_{b_{5},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.21)
$\mathcal{F}_{b_{6}}$ $=\mathcal{F}_{b_{4}},$ (3.22)
$\mathcal{F}_{b_{7}}$ $=\mathcal{F}_{b_{5}},$ (3.23)
where
$\mathcal{I}_{b_{i},K^{\prime}}$ $=\mathcal{J}_{b_{i}}(k_{1},k_{2},k_{3}),$ (3.24)
$\mathcal{I}_{b_{i},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(-k_{1},k_{2},k_{3}),$ (3.25)
$\mathcal{I}_{b_{i},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{i}}(k_{1},-k_{2},k_{3}),$ (3.26)
$\mathcal{I}_{b_{i},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{i}}(k_{1},k_{2},-k_{3}),$ (3.27)
with
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}},$ (3.28)
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}),$ (3.29)
$\mathcal{J}_{b_{2}}(k_{1},k_{2},k_{3})$ $:=-\frac{1}{K^{\prime 2}}\biggl{[}c_{s}^{3}k_{1}^{3}+2c_{s}^{2}c_{h}k_{1}^{2}(k_{2}+k_{3})$
$\qquad\qquad\quad+2c_{s}c_{h}^{2}k_{1}(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})+c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})\biggr{]},$ (3.30)
$\mathcal{S}_{b_{2}}(k_{1},k_{2},k_{3})$ $:=e^{iK^{\prime}\eta_{0}}\biggl{[}\frac{c_{h}}{K^{\prime 2}}(c_{h}^{2}k_{2}k_{3}(k_{2}+k_{3})+c_{s}^{2}k_{1}^{2}(k_{2}+k_{3})$
$\qquad\qquad\quad+c_{s}c_{h}k_{1}(k_{2}^{2}+4k_{2}k_{3}+k_{3}^{2}))+\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr{]},$ (3.31)
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}},$ (3.32)
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}),$ (3.33)
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=\frac{2}{K^{\prime 3}},$ (3.34)
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2}),$ (3.35)
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3})$ $:=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}},$ (3.36)
$\mathcal{S}_{b_{5}}(k_{1},k_{2},k_{3})$ $:=e^{iK^{\prime}\eta_{0}}\biggl{\{}\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}$
$\qquad\qquad\quad+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr{\}}.$ (3.37)
For the Bunch-Davies state limit ($\beta^{(s)}_{k}\to 0$), this expression gives the previous result in Ref. [101]. In particular, in the squeezed limit the difference from the case with the Bunch-Davies state appears in $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}$, and $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}/\mathcal{I}_{b_{i},K^{\prime}}$ corresponds to the characteristic scale-dependent enhancement of the squeezed bispectrum with the non-Bunch-Davies states. By taking the squeezed limit ($k_{L}=k_{1}\ll k_{2}\simeq k_{3}\simeq k_{S}$), we have
$K^{\prime}\simeq K_{1}^{\prime}\simeq 2c_{h}k_{S},K_{2}^{\prime}\simeq K_{3}^{\prime}\simeq c_{s}k_{L},$ (3.38)
and $\mathcal{I}_{b_{i},\tilde{K}}$ can be simply estimated as
$\mathcal{I}_{b_{1},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{1},K_{1}^{\prime}}\simeq\frac{1}{2c_{h}k_{S}},\ \mathcal{I}_{b_{1},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{1},K_{3}^{\prime}}\simeq i\eta_{0},$ (3.39)
$\mathcal{I}_{b_{2},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{2},K_{1}^{\prime}}\simeq\frac{3c_{h}k_{S}}{2},\ \mathcal{I}_{b_{2},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{2},K_{3}^{\prime}}\simeq-ic_{h}^{2}k_{S}^{2}\eta_{0},$ (3.40)
$\mathcal{I}_{b_{3},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{3},K_{1}^{\prime}}\simeq\frac{3}{4c_{h}k_{S}},\ \mathcal{I}_{b_{3},K_{2}^{\prime}}\simeq-\mathcal{I}_{b_{3},K_{3}^{\prime}}\simeq\frac{c_{h}k_{S}}{2}\eta_{0}^{2},$ (3.41)
$\mathcal{I}_{b_{4},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{4},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{4},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{4},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3},$ (3.42)
$\mathcal{I}_{b_{5},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{5},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{5},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{5},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3}.$ (3.43)
Accordingly, we have shown explicitly that the STT bispectrum with the non-Bunch-Davies states in the squeezed limit can be enhanced in proportion to powers of $k_{S}\eta_{0}$, compared to that with the Bunch-Davies state. In Sec. 4, we will investigate how those enhancements can leave their imprints on the SGWB anisotropies.
4 SGWB anisotropies
4.1 General expression for the anisotropies
It has been shown in Refs. [102, 103, 58, 104, 105, 21, 22, 23, 24, 26, 27, 28, 29, 30, 32] that the primordial non-Gaussianities result in spatial modulations of the primordial scalar and tensor power spectra. In particular, the STT bispectrum induces the directional dependence in the tensor power spectrum: [27, 29, 32]
$\mathcal{P}_{h}^{(s)}(k,\hat{n})=\bar{\mathcal{P}}^{(s)}_{h}(k)\biggl{[}1+\sum_{s^{\prime}}\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q})\biggr{]},$ (4.1)
where $\hat{n}:=-{\bf k}/k$ and $d:=\eta_{\rm now}-\eta_{\rm ent}$ with $\eta_{\rm now}$ and $\eta_{\rm ent}$ being the conformal time at present and at the horizon reentry of inflationary GWs for each $k$ mode, respectively. Since we consider the nHz or deci-Hz inflationary GWs that re-entered the horizon long before the present time, that is, $\eta_{\rm now}\gg\eta_{\rm ent}$, we assume $d\simeq\eta_{\rm now}$. Hereinafter, a bar means the homogeneous component. In the above expression, $F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})$ can be characterized by the cross bispectrum (3.7) as
$F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q}):=\mathcal{B}^{ss^{\prime}}_{shh}({\bf k}-{\bf q}/2,-{\bf k}-{\bf q}/2,{\bf q})|_{q\to 0}\biggl{(}\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h}(k)\biggr{)}^{-1}\biggl{(}\frac{2\pi^{2}}{q^{3}}\mathcal{P}_{\zeta}(q)\biggr{)}^{-1},$ (4.2)
with
${\bf k}_{1}={\bf q},~{}{\bf k}_{2}={\bf k}-{\bf q}/2,~{}{\bf k}_{3}=-{\bf k}-{\bf q}/2.$ (4.3)
Note that $k$ and $q$ in this section corresponds to $k_{L}(=k_{1})$ and $k_{S}(\simeq k_{2}\simeq k_{3})$ used in the above sections, respectively. As for the SGWB, the density parameter per logarithmic wavenumber:
$\Omega_{\rm GW}:=\frac{1}{\rho_{\rm cr}}\frac{{\rm d}\rho_{\rm GW}}{{\rm d}\ln k},$ (4.4)
has been widely used, where $\rho_{\rm cr}$ is the critical density at present, and for the inflationary GWs it is proportional to the primordial tensor power spectrum. Thus, the anisotropies of the SGWB induced from the primordial STT bispectrum can be characterized as [27, 29, 32]
$\Omega_{\rm GW}(k,\hat{n})=\bar{\Omega}_{\rm GW}(k)\biggl{[}1+\delta_{\rm GW}(k,\hat{n})\biggr{]},$ (4.5)
where
$\delta_{\rm GW}(k,\hat{n}):=\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}\sum_{ss^{\prime}}\tilde{F}^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q}),$ (4.6)
with
$\tilde{F}_{\rm NL}^{ss^{\prime}}({\bf k},{\bf q}):=\frac{\bar{\mathcal{P}}^{(s)}_{h}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})}{\sum_{s}\bar{\mathcal{P}}^{(s)}_{h}}.$ (4.7)
One can expand the anisotropies in terms of the spherical harmonics, the expansion coefficients of which are of the form
$\delta^{\rm GW}_{\ell m}(k)=\int{\rm d}^{2}\hat{n}Y^{*}_{\ell m}(\hat{n})\delta_{\rm GW}(k,\hat{n}).$ (4.8)
In general, the three-point correlation functions are proportional to products of the polarization tensors that include angular dependence and $F^{s_{2}s_{3}}_{\rm NL}$ can be decomposed in terms of the spherical harmonics as
$\tilde{F}^{s_{2}s_{3}}_{\rm NL}({\bf k},{\bf q})=Y_{00}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,0}+Y_{20}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,2}+Y_{40}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,4}+\cdots=\sum_{L={\rm even}}f^{s_{2}s_{3}}_{{\rm NL},L}Y_{L0}(\hat{n}\cdot\hat{q}),$ (4.9)
where $\hat{q}:={\bf q}/q$. With this parametrization, we have
$\delta^{\rm GW}_{\ell m}(k)$ $=4\pi\sum_{L^{\prime},M,M^{\prime}}\sum_{L={\rm even}}i^{L^{\prime}}\sqrt{(2L^{\prime}+1)(2\ell+1)}\begin{pmatrix}\ell&L^{\prime}&L\\ 0&0&0\end{pmatrix}\begin{pmatrix}\ell&L^{\prime}&L\\ m&M^{\prime}&M\end{pmatrix}$
×$\quad\times\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}Y_{L^{\prime}M^{\prime}}(\hat{q})Y_{LM}(\hat{q})j_{L^{\prime}}(qd)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)\zeta_{\bf q},$ (4.10)
with $\begin{pmatrix}l_{1}&l_{2}&l_{3}\\ L_{1}&L_{2}&L_{3}\end{pmatrix}$ and $j_{L}(x)$ being the Wigner-$3j$ symbol and spherical Bessel function respectively.
4.2 Angular correlation for the SGWB anisotropies
Based on the previous expression, as statistical observables for the SGWB anisotropies, we can introduce the angular auto-correlation function $C^{\rm GW}_{\ell}$ as
$C^{\rm GW}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)\delta^{\rm GW*}_{\ell m}(k)\rangle=\sum_{L={\rm even}}C^{\rm GW}_{\ell,L}(k),$ (4.11)
where
$C^{\rm GW}_{\ell,L}(k)$ $=(2L+1)\sum_{J,J^{\prime}}i^{J-J^{\prime}}(2J+1)(2J^{\prime}+1)\biggl{[}\begin{pmatrix}\ell&J&L\\ 0&0&0\end{pmatrix}\biggr{]}^{2}\biggl{[}\begin{pmatrix}\ell&J^{\prime}&L\\ 0&0&0\end{pmatrix}\biggr{]}^{2}$
×$\quad\times\int\frac{{\rm d}q}{q}j_{J}(qd)j_{J^{\prime}}(qd)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3},s^{\prime}_{2},s^{\prime}_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)f^{s^{\prime}_{2}s^{\prime}_{3}}_{{\rm NL},L}(k,q).$ (4.12)
We also consider the cross-correlation between the SGWB and CMB anisotropies. The CMB anisotropies can be expanded in terms of the spherical harmonics as
$X(\hat{n})=\sum_{\ell m}a^{X}_{\ell m}Y^{*}_{\ell m}(\hat{n}),$ (4.13)
with
$a^{X}_{\ell m}=4\pi i^{\ell}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\mathcal{T}^{X}_{\ell}(p)Y^{*}_{\ell m}(\hat{p})\zeta({\bf p}),$ (4.14)
where the $X$ stands for $T$ (temperature fluctuations) or $E$ (E-mode polarization). [^3] As has been studied in Ref. [32], the cross-correlation between the B-mode polarization and the SGWB anisotropy sourced by the long-wavelength tensor mode exists in parity-violating gravitational theories. In the same way as the angular auto-correlation function of the SGWB anisotropies, the cross-correlation function of the CMB and SGWB anisotropies can be expressed as
$C^{\rm GW-X}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)a^{X*}_{\ell m}\rangle=\sum_{L={\rm even}}C^{\rm GW-X}_{\ell,L}(k),$ (4.15)
with
$C^{\rm GW-X}_{\ell,L}(k)=4\pi\sqrt{\frac{2L+1}{4\pi}}\sum_{J}i^{J-\ell}(2J+1)\biggl{[}\begin{pmatrix}\ell&J&L\\ 0&0&0\end{pmatrix}\biggr{]}^{2}\int\frac{{\rm d}q}{q}j_{J}(qd)\mathcal{T}^{X}_{\ell}(q)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q),$ (4.16)
where $\mathcal{T}^{X}_{\ell}(q)$ is the transfer function for the CMB anisotropies.
4.3 SGWB anisotropies from inflation with the non-Bunch-Davies states
Then, let us evaluate the angular correlation function sourced from the primordial scalar-tensor-tensor bispectrum with the non-Bunch-Davies states. In the present setup, one can take the squeezed limit to the STT bispectrum (3.10) straightforwardly as
$\frac{\mathcal{B}_{shh}^{ss}}{P_{s}(q)P_{h}^{(s)}(k)}$ $=4\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}\right){\rm Re}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
$\quad-8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right){\rm Im}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]c_{h}k\eta_{0}$
$\quad+\frac{8b_{2}}{\mathcal{F}_{T}}{\rm Im}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]\frac{1}{c_{h}k\eta_{0}}$
$\quad-\frac{4c_{s}c_{h}b_{3}}{\mathcal{F}_{T}}\,{\rm Re}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times(\hat{n}\cdot\hat{q})\times c_{h}k\eta_{0}\times c_{s}q\eta_{0}+O((q/k)^{2}).$ (4.17)
Note that here we assume $s_{2}=s_{3}=s$. This is because as can be seen in the expressions for $\mathcal{V}_{i}$ at the squeezed limit studied in Appendix C, $\mathcal{V}_{i}$ are suppressed at most in the order of $q/k$ for the case with $s_{2}=-s_{3}$. Thus, in the leading order, not only the power spectrum but also the bispectrum depend on the choice of the helicity only through the helicity dependence of the Bogoliubov coefficients. In the following subsections, we will investigate the correlation functions of the SGWB anisotropies for the two cases of the Bogoliubov coefficients which are introduced in section 2.2.
4.3.1 Peaked spectrum
The fourth line in the squeezed bispectrum  (4.17), which is proportional to $\hat{n}\cdot\hat{q}$, has
$(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}}).$ (4.18)
By substituting the assumptions for the Bogoliubov coefficients given by Eqs. (2.31) and (2.32), one can find that it is reduced to ${\rm Re}[\beta^{(s)}_{k_{2}}]-{\rm Re}[\beta^{(s)}_{k_{3}}]$ and at the leading order in the squeezed limit this should be the order of $(\hat{n}\cdot\hat{q})q$. Thus, in the squeezed limit the fourth line of Eq. (4.17) is found to be proportional to $(\hat{n}\cdot\hat{q})^{2}\times(c_{s}q\eta_{0})^{2}$. Since we focus on the phase after the long-wavelength mode $q$ crosses the sound horizon, i.e., the phase characterized by $|c_{s}q\eta_{0}|\ll 1$, there is no enhancement for the term proportional to $(\hat{n}\cdot\hat{q})^{2}$ in the bispectrum. Accordingly, it is enough to consider the terms proportional to $(\hat{n}\cdot\hat{q})^{0}$.
From Eqs. (2.31) and (2.32), up to the leading order in $q/k$, we can also find
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]\approx 1,$ (4.19)
${\rm Im}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
$\qquad\qquad\qquad\approx{\rm Im}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
$\qquad\qquad\qquad\approx\frac{2A^{2}}{2A^{2}-2Ae^{\ln^{2}(k/k_{*})/\Delta^{2}}+e^{2\ln^{2}(k/k_{*})/\Delta^{2}}}=:\mathcal{A}(k,k_{*}).$ (4.20)
For $k\approx k_{*}$ and $A\gg 1$, $\mathcal{A}(k,k_{*})\approx 1$. As a result, the non-linearity parameter in the squeezed limit of the scalar-tensor-tensor bispectrum $F_{\rm NL}^{s}$ for the peaked tensor power spectrum can be rewritten as
$\tilde{F}_{\rm NL}^{ss}:=\frac{\mathcal{B}^{ss}_{shh}|_{q\to 0}}{P_{s}(q)P_{h}^{(s)}(k)}$ $=\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*}),$ (4.21)
and also
$f^{ss}_{{\rm NL},0}(k)=\frac{\tilde{F}^{ss}_{\rm NL}}{Y_{00}}=2\sqrt{\pi}\biggl{\{}\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*})\biggr{\}}.$ (4.22)
Note that the leading-order non-linearity parameter is independent of $q$ in the squeeze limit.
We then compute the correlation functions of the SGWB anisotropies. From Eq. (4.12), we obtain
$C^{\rm GW}_{\ell}(k)=\frac{4\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)^{2}}{2\ell(\ell+1)},$ (4.23)
where we assumed the scale-invariant scalar power spectrum and we used
$\int^{\infty}_{0}\frac{{\rm d}x}{x}j_{\ell}(x)j_{\ell}(x)=\frac{1}{2\ell(\ell+1)}.$ (4.24)
For the cross-correlations between the SGWB and CMB anisotropies, we only consider the temperature fluctuation since the amplitude of the E-mode polarization is much smaller than that of the temperature fluctuation. Since we are interested in the temperature fluctuation on a large scale where the SW effect dominates other effects, as the transfer function for the temperature anisotropy we use the following approximate form:
$\mathcal{T}^{T}_{\ell}(k)\simeq\frac{1}{5}j_{\ell}(k(\eta_{\rm now}-\eta_{\rm dec})),$ (4.25)
where $\eta_{\rm dec}$ denotes the decoupling time. In the calculation of the correlation functions, we use $\eta_{\rm dec}/\eta_{\rm now}\sim 0.02$. As a result, we can rewrite $C_{\ell}^{\rm GW-T}$ as
$C^{\rm GW-T}_{\ell}(k)\simeq\frac{4\sqrt{\pi}}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\int\frac{{\rm d}q}{q}j_{\ell}(qd)j_{\ell}(q(\eta_{\rm now}-\eta_{\rm dec})),$ (4.26)
the integral of which can be performed analytically as
$\int_{0}^{\infty}\frac{{\rm d}q}{q}j_{\ell}(qd)j_{\ell}(q(\eta_{\rm now}-\eta_{\rm dec}))$ $=\int_{0}^{\infty}\frac{{\rm d}(qd)}{qd}j_{\ell}(qd)j_{\ell}(qd(\eta_{\rm now}-\eta_{\rm dec})/d)$
$=\frac{1}{4}c^{\ell}\sqrt{\pi}\Gamma(\ell)\frac{{}_{2}F_{1}\biggl{(}-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr{)}}{\Gamma(\frac{3}{2}+\ell)},$ (4.27)
where ${}_{2}F_{1}(a,b,c;z)$ is the hypergeometrical function, and we introduced
$c:=\frac{\eta_{\rm now}-\eta_{\rm dec}}{d}\simeq\frac{\eta_{\rm now}-\eta_{\rm dec}}{\eta_{\rm now}}.$ (4.28)
Note that by taking $c\to 1$ to Eq. (4.27), we obtain the same result with Eq. (4.24). The resultant form of $C_{\ell}^{{\rm GW}-T}$ reads
$C_{\ell}^{\rm GW-T}(k)\simeq\frac{\pi}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\times c^{\ell}\frac{\Gamma(\ell)}{\Gamma(\ell+\frac{3}{2})}{}_{2}F_{1}\biggl{(}-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr{)}.$ (4.29)
The plots of normalized $C_{\ell}$s are shown in Fig. 1.
Refer to caption Refer to caption
Figure 1: Left: The plot of $C_{\ell}^{\rm GW}$ normalized by $\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)^{2}$. Right: The plot of $C_{\ell}^{\rm GW-T}$ normalized by $\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)$.
Let us investigate the typical amplitudes of the SGWB anisotropies induced by the primordial STT bispectrum with the non-Bunch-Davies vacuum states. To do so, we first need to inspect two theoretical constraints on $b_{i}$ and $\eta_{0}$. The first constraint comes from an argument on the validity of the linear perturbation theory. In the present paper, we discuss this point by requiring that the STT interaction is smaller than the quadratic tensor interaction. By requiring $\mathcal{L}^{(2)}_{h}\geq\mathcal{L}^{(3)}_{shh}$ where $\mathcal{L}^{(2)}_{h}$ stands for the quadratic Lagrangian of the tensor perturbations, we have
$\frac{b_{1}}{\mathcal{G}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1},$ (4.30)
$\frac{b_{2}}{\mathcal{F}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1}.$ (4.31)
In light of that the curvature perturbation is on the superhorizon scales at $\eta=\eta_{0}$, we can use the current constraint on the amplitude of $\zeta$ as $|\zeta(\eta_{0})|\sim 10^{-5}$ [106]. Another constraint is obtained from an argument on backreaction from the excited tensor modes. Once the modes get excited from the Bunch-Davies state, the excited modes cause backreaction to the inflationary background [107]. In the Horndeski theory, the excited tensor mode has the following energy density [67]:
$\rho_{\rm excited}=\frac{c_{h}}{a^{4}(\eta)}\int|\beta^{(s)}_{k}|^{2}k^{3}{\rm d}k.$ (4.32)
In order for the backreaction from the excited modes not to spoil the background dynamics of the slow-roll inflation, we compare this to the homogeneous part of the Friedmann equation $\mathcal{E}_{b}$. In the standard slow-roll inflation, the homogeneous part corresponds to the energy density of inflaton as $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$. In the general inflation model considered here, we similarly assume $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$ for simplicity. Then, the backreaction constraint, i.e., $\rho_{\rm excited}\lesssim\mathcal{E}_{b}$, reads
$\frac{1}{a(\eta_{0})^{4}}A^{2}e^{2\Delta^{2}}k_{*}^{4}\Delta\lesssim M_{\rm pl}^{2}H_{\rm inf}^{2}.$ (4.33)
If we assume $A=-3000$, $\Delta=0.45$ and $H_{\rm inf}=10^{13}~{}{\rm GeV}$ to explain the recent PTA results, the above reads
$-k_{*}\eta_{0}\lesssim 10.$ (4.34)
From Eqs. (4.30), (4.31), and (4.34), we can estimate the theoretical upper bound on $f_{\rm NL}$ as
$f^{ss}_{{\rm NL},0}(k_{*})\lesssim 10^{7}.$ (4.35)
We should also discuss an observational constraint on $f^{ss}_{{\rm NL},0}(k)$. So far, the CMB temperature bispectrum $B_{TTT}$ (i.e., a non-linearity parameter for a scalar non-Gaussianity) has been roughly constrained as $B_{TTT}/P_{\zeta}^{2}\lesssim 10$ [108]. Since both scalar and tensor perturbations generate the CMB temperature fluctuations (i.e., $\mathcal{B}_{shh}$ contributes to $B_{TTT}$), the above constraint naively applies to the STT bispectrum as
$\mathcal{B}_{shh}|_{k_{\rm CMB}}\sim\mathcal{O}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr{)}\times P_{\zeta}|_{k_{\rm CMB}}P_{h}|_{k_{\rm CMB}}\lesssim 10P^{2}_{\zeta}|_{k_{\rm CMB}},$ (4.36)
which yields the following constraint
$\mathcal{O}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr{)}\lesssim 10^{3}\,\left(\frac{r|_{k_{\rm CMB}}}{10^{-2}}\right)^{-1}.$ (4.37)
For $r|_{k_{\rm CMB}}=10^{-2}$, we obtain a stronger constraint on $f^{ss}_{{\rm NL},0}(k)$ (i.e., $f^{ss}_{{\rm NL},0}(k)\lesssim 10^{3}$) than the theoretical one in Eq. (4.35). Finally, we can estimate the possible maximum values of $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$. Since we have assumed the Bunch-Davies scalar mode, we have $\mathcal{P}_{\zeta}\simeq 2\times 10^{-9}$. Also, the upper bound on $f^{ss}_{{\rm NL},0}(k)$ is evaluated as $10^{3}$ as discussed above. As a result, we can find that $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$ can reach $10^{-4}$ and $10^{-7}$, respectively. The allowed values of the correlation functions are shown in Fig. 2.
Refer to caption Refer to caption
Figure 2: Left: The allowed values of $C_{\ell}^{\rm GW}$. Right: The allowed values of $C_{\ell}^{\rm GW-T}$.
Let us mention the possibility of obtaining a large value of $b_{1,2}$ in the framework of slow-roll inflation. So far, we have shown that $f^{ss}_{{\rm NL},0}(k)$ can be enhanced by a factor $-k\eta_{0}\,(\lesssim 10)$ due to the consequence of the non-Bunch-Davies effects. The largest enhancement of $f^{ss}_{{\rm NL},0}(k)$ can thus arise from large values of the coupling functions of the cubic interactions $b_{1,2}$. Thus, one might expect that similar enhancements could be predicted from inflation with the Bunch-Davies state. However, in that case, the squeezed bispectrum is expected to satisfy the so-called consistency relation between the spectral index $n_{t}$ and the amplitude of the squeezed bispectrum [37, 59]. To show it explicitly in the Horndeski theory, one would need to take into account slow-roll corrections to the scale factor and the mode functions as has been studied in the context of the scalar non-Gaussianity in Ref. [109]. Without taking those into account, one can show at least that the squeezed bispectrum with the Bunch-Davies state is suppressed by the slow-roll parameters. Now the terms in the case of the Bunch-Davies state correspond to the ones irrelevant with $k\eta_{0}$ in Eq. (4.22). The explicit forms of $b_{1}$ and $b_{2}$ are obtained as
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl{\{}1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl{[}-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}\biggr{]}\biggr{\}},$ (4.38)
$b_{2}$ $=\frac{1}{8}\frac{H\mathcal{G}_{T}^{2}}{\Theta}+\frac{1}{8}\frac{\mathcal{G}_{T}}{\Theta}\dot{\mathcal{G}}_{T}+\frac{\mathcal{G}_{T}}{8}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}-\frac{\mathcal{F}_{T}}{8}.$ (4.39)
The terms without the time derivatives cancel with each other only in the following combination:
$\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}$ $=\frac{\sigma}{8c_{h}^{2}}\biggl{(}2\epsilon-f_{T}+3g_{T}+\frac{2\dot{\sigma}}{H}\biggr{)},$ (4.40)
with $\sigma:=H\mathcal{G}_{T}/\Theta$. Note that $\sigma=1$ in the k-essence theory. The above is nothing but the combination appearing in the case of the Bunch-Davies state and is suppressed by the slow-roll parameters. However, for different combinations (e.g., $b_{1}/\mathcal{G}_{T}+b_{2}/\mathcal{F}_{T}$), there are terms that are not proportional to the slow-roll parameters. To see more details about the enhancement of $f^{ss}_{{\rm NL},0}(k)$, let us express our $f^{ss}_{{\rm NL},0}(k)$ in terms of $f_{\rm NL,BD}$ which is the result in the case of the Bunch-Davies state as
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl{(}f_{\rm NL,BD}-\sqrt{\pi}\frac{b_{2}}{\mathcal{F}_{T}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr{)},$ (4.41)
where
$f_{\rm NL,BD}:=8\sqrt{\pi}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}\biggr{)}.$ (4.42)
If both scalar and tensor power spectra are almost scale invariant on the CMB scales, we have
$r|_{k_{\rm CMB}}\simeq 16\frac{\mathcal{F}_{S}}{\mathcal{F}_{T}}\frac{c_{s}}{c_{h}}.$ (4.43)
Taking this into account and using $b_{2}=\mathcal{F}_{S}/8$, we can rewrite $f^{ss}_{{\rm NL},0}(k)$ as
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl{(}f_{\rm NL,BD}-\frac{\sqrt{\pi}}{128}\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr{)}.$ (4.44)
We can thus estimate the deviation of $f_{\rm NL}$ from $f_{\rm NL,BD}$ as
$|f^{ss}_{{\rm NL},0}(k)-f_{\rm NL,BD}|\leq\mathcal{O}\biggl{(}\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\biggr{)}\leq\mathcal{O}\biggl{(}10^{-2}\frac{c_{h}}{c_{s}}\biggr{)},$ (4.45)
where we assumed $r|_{k_{\rm CMB}}\sim 10^{-2}$ in the last step, which indicates that one needs a small $c_{s}$. This is dangerous at least within the k-essence theory where a small $c_{s}$ always indicates a large scalar non-Gaussianity [38]. However, this is not the case in more general models of inflation, and in the Horndeski theory, a small $c_{s}$ does not necessarily mean a large scalar non-Gaussianity. Therefore, we conclude that inflation with the non-Bunch-Davies states can potentially enhance $f^{ss}_{{\rm NL},0}(k)$ up to $10^{3}$ and the SGWB anisotropies $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ up to $10^{-4}$ and $10^{-7}$, respectively.
Before moving to the scale-invariant case, let us compare the amplitudes of the SGWB anisotropies from inflation with the non-Bunch-Davies state to those originating from the propagation of the GWs at a late time. On large scales, the dominant contributions to the anisotropies originate from the SW effect as in the CMB fluctuations, the auto- and cross-correlation functions originating from which are, respectively, of the form, [29]
$C_{\ell}^{\rm GW,SW}$ $=\frac{128\pi}{9\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 3\times 10^{-9}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right),$ (4.46)
$C_{\ell}^{\rm GW-T,SW}$ $=\frac{16\pi}{15\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-10}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (4.47)
Due to the enhancements by $f^{ss}_{{\rm NL},0}(k)$, the anisotropies originating from inflation with the non-Bunch-Davies state can dominate those from the SW effect.
4.3.2 Nearly-scale-invariant spectrum
In the case of the $k$-independent Bogoliubov coefficients, [^4] In the scale-invariant case, the contribution which is proportional to $\hat{n}\cdot\hat{q}$ in the scalar-tensor-tensor bispectrum is zero in the squeezed limit. we have
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]=1,$ (4.48)
${\rm Im}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
$\qquad\qquad\qquad={\rm Im}\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}+\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]$
$\qquad\qquad\qquad=-2B.$ (4.49)
Then, $f^{ss}_{{\rm NL},0}(k)$ reads
$f^{ss}_{{\rm NL},0}(k)=2\sqrt{\pi}\biggl{\{}\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)+2B\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\biggr{\}}.$ (4.50)
Since the above $f^{ss}_{{\rm NL},0}(k)$ is independent of $q$, calculations of the correlation functions can be similarly performed as in the previous case with peaked power spectrum, and $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ are given by Eqs. (4.23) and (4.29), respectively.
As for the constraints on the non-linearity parameters, the constraints on the coupling functions of the cubic interactions $b_{1,2}$ are also expected to be roughly given by Eqs. (4.30), (4.31) and (4.37). On the other hand, the theoretical backreaction constraint reads
$c_{h}B^{2}M_{\rm Pl}^{4}\lesssim M_{\rm Pl}^{2}H_{\rm inf}^{2},$ (4.51)
where we have introduced a cut-off scale $\Lambda=M_{\rm Pl}$ for the physical momentum ($k/a(\eta)$) at $\eta=\eta_{0}$. We thus have $B\lesssim 10^{-5}$ for $H_{\rm inf}=10^{13}~{}{\rm GeV}$ and $c_{h}\simeq 1$, and
$-k\eta_{0}=\frac{k}{a(\eta_{0})H_{\rm inf}}=\frac{k_{\rm max}}{a(\eta_{0})H_{\rm inf}}\frac{k}{k_{\rm max}}=\frac{M_{\rm Pl}}{H_{\rm inf}}\frac{k}{k_{\rm max}}\sim 10^{5}\frac{k}{k_{\rm max}}\leq 10^{5},$ (4.52)
which indicates
$(-k\eta_{0})B\lesssim 1.$ (4.53)
Here, one has $(-k\eta_{0})B\sim 1$ only for the case of $k_{\rm max}=k$, i.e., $k/a(\eta_{0})\sim M_{\rm Pl}$. Thus, in the scale-invariant case the amplitude of SGWB anisotropies is found to be determined by the cubic coupling functions, $b_{1,2}$ even with taking the non-Bunch-Davies vacuum states into account. Nevertheless, contrary to the Bunch-Davies vacuum, we can expect the violation of Maldacena’s consistency relation as discussed above, and hence the maximum amplitudes of the anisotropies in the scale-invariant spectrum case are the same as those in the peaked spectrum case.
4.4 Detectability of the SGWB anisotropies
Based on the formulae for the correlation functions of the SGWB anisotropies induced by the inflationary STT bispectrum with the non-Bunch-Davies states shown in the above sections, let us discuss the detectability of such inflationary SGWB anisotropies in future experiments. The detectability of the SGWB anisotropies has been discussed for space-based interferometers BBO/DECIGO and Square Kilometre Array (SKA) as a PTA experiment in Ref. [29], where a noise angular spectrum $N_{\ell}$ for each experiment has been computed for the case of a blue-tilted tensor spectrum. In fact, in our paper, we have considered the peaked tensor power spectrum with the non-Bunch-Davies state to explain the recent SGWB data reported by the PTA experiments such as NANOGrav, and it is not exactly the same as the one assumed in Ref. [29]. However, as discussed in the previous section 2.2.1, by setting the model parameters appropriately, our model can reproduce the blue-tilted tensor power spectrum for the relevant frequency range of the PTA experiments. Thus, we employ their result of the noise spectrum when discussing the detectability with future PTA experiments such as SKA. For the scale-invariant case discussed in Sec. 4.3.2, we employ the noise spectrum computed in Ref. [110]. We investigate the detectability of the SGWB anisotropies for the case of the scale-independent and -dependent Bogoliubov coefficients with BBO/DECIGO and SKA, respectively.
Since the power spectra of the SGWB anisotropies induced from the primordial STT bispectrum with the non-Bunch-Davies states decrease as the multipole moment $\ell$ increases as shown in the previous section, and also the amplitudes of the noise spectra rapidly increase at $\ell>10$ [29, 110], in the estimation of the detectability we only consider the $-\ell$ region as $\ell\leq 10$. Following Refs. [29, 110], for $\ell\leq 10$, the typical amplitudes of the noise spectrum are given as $N_{\ell}=\mathcal{O}(10^{-30})$ in the case of the scale-invariant spectrum for BBO/DECIGO and $N_{\ell}=\mathcal{O}(10^{-26})$ in the case of the blue-tilted spectrum for SKA, respectively. The noise used for computing the signal-to-noise ratio (SNR) of the SGWB anisotropies is
$N_{\ell}^{\rm GW}:=\frac{N_{\ell}}{\bar{\Omega}^{2}_{\rm GW}}.$ (4.54)
For the auto-correlation function, we inspect $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ as the SNR, [^5] Exactly speaking, the SNR squared is evaluated by computing [10, 32] $\sum_{\ell}T_{\rm obs}\int{\rm d}f(2\ell+1)\frac{C_{\ell}^{\rm GW}}{N_{\ell}^{\rm GW}},$ (4.55) where $T_{\rm obs}$ is the total time for observations. In the present paper, we use $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ for the purpose of order estimation. that is, the threshold for the detection is roughly estimated as $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}\gtrsim 1$. In our case, $C_{\ell}^{\rm GW}$ is at most of $\mathcal{O}(10^{-4})$ for both scale-independent and peaked Bogoliubov coefficients. For the DECIGO/BBO case, assuming the scale-invariant power spectrum, we have $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-16})$ for $r=10^{-2}$ and hence $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}\ll 1$. For the PTA case, assuming that the amplitude of the SGWB is that of reported signals with recent PTA experiments, that is, $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$ at nHz, we can expect $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}<O(10^{4})$. Thus, if the reported nHz SGWB is sourced by the inflationary GWs with non-Bunch-Davies state, in future PTA experiments such as SKA it could be possible to test the vacuum state during inflation by the observation of the SGWB anisotropies.
As for the scale-invariant case with DECIGO/BBO experiments, let us briefly investigate the detectability by making use of the cross-correlation between GWs and CMB. In this case, as we have discussed, $C^{\rm GW}_{\ell}\ll N^{\rm GW}_{\ell}$, and then the SNR for the cross-correlation can be roughly computed as [29]  [^6] For the case of $N^{\rm GW}_{\ell}\lesssim C^{\rm GW}_{\ell}$, the SNR can be roughly given by [29] ${\rm SNR}\simeq\frac{C^{\rm GW-T}_{\ell}}{\sqrt{(C^{\rm GW}_{\ell}+N^{\rm GW}_{\ell})C^{\rm TT}_{\ell}}}.$ (4.56) For the SKA case (i.e., $N_{\ell}=\mathcal{O}(10^{-26})$), by assuming $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$, the above SNR can reach unity for $f^{ss}_{{\rm NL},0}\geq\mathcal{O}(1)$. Therefore, in light of the upper bound on $f^{ss}_{{\rm NL},0}$, the amplitudes of the SGWB anisotropies for the peaked-spectrum case can be large enough to be tested by both auto- and cross-correlation.
${\rm SNR}\simeq\frac{C_{\ell}^{\rm GW-T}}{\sqrt{N_{\ell}^{\rm GW}C_{\ell}^{\rm TT}}},$ (4.57)
where $C_{\ell}^{\rm TT}$ is the angular power spectrum of the CMB temperature fluctuation:
$C^{\rm TT}_{\ell}=\frac{1}{2\ell+1}\sum_{m}\langle a^{T}_{\ell m}a^{T*}_{\ell m}\rangle\simeq\frac{2\pi}{25\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-11}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (4.58)
Here, we have used the approximated form of the transfer function of the temperature fluctuation in Eq. (4.25). Thus, even if the cross-correlation would be used for the DECIGO/BBO experiments, it is hard to detect the SGWB anisotropies induced from the inflationary STT bispectrum with the scale-independent Bogoliubov coefficient. Note that, similarly to the SKA case, one can consider the tensor spectrum peaked at the frequency band targeted by DECIGO/BBO to enhance $\bar{\Omega}_{\rm GW}$ without contradicting CMB experiments by employing the scale-dependent Bogoliubov coefficient. For such a case, we need to estimate the noise spectrum with the peaked tensor power spectrum for the DECIGO/BBO experiments and also investigate the backreaction problem carefully. We will leave the detailed analysis of these points to our future work.
5 Summary
The SGWB anisotropies have recently attracted a lot of attention as a tool to distinguish the source of SGWBs and a probe of the new physics in the early universe. In this paper, we have investigated the SGWB anisotropies induced by the primordial non-Gaussianity during inflation. Although it has been mentioned that the characteristic primordial non-Gaussianity with non-Bunch-Davies states, which brings the coupling between short- and long-wavelength modes, can induce the spatial anisotropies on SGWB, precise theoretical predictions were not made in previous studies. In this paper, we have explicitly formulated the primordial scalar-tensor-tensor (STT) bispectrum induced during inflation in the context of Horndeski theory with non-Bunch-Davies states and investigated the STT bispectrum in the squeezed limit, which characterizes the coupling between long-wavelength scalar mode and short-wavelength tensor modes. In the formulation, we carefully clarified the Trans-Planckian problem for the short-wavelength tensor modes in the non-Bunch-Davies state and then derived the condition to avoid it. In the CMB observations, due to the relatively small hierarchy between the longest and shortest modes which are in the observational range (roughly estimated as $k_{\rm S}/k_{\rm L}=O(10^{3})$), the aforementioned Trans-Planckian problem is not so serious. On the other hand, for the SGWB anisotropies, a large hierarchy between long- and short-wavelength modes (roughly estimated as $k_{\rm S}/k_{\rm L}=O(10^{8})$ (for PTA experiments), and $O(10^{16})$ (for DECIGO/BBO experiments)) is required, and then close attention must be paid to this problem. We found that by avoiding the Trans-Planckian problem the strong scale-dependent enhancement of the STT bispectrum in the squeezed limit in the case with the non-Bunch-Davies states, which was considered in the previous works, can not be expected. Instead, another scale dependence appears in the STT bispectrum which also depends on the non-Bunch-Davies excitation time.
To obtain the theoretical predictions for the SGWB anisotropies, with future GW observations in mind, we assumed the two configurations of the Bogoliubov coefficients: one is for the highly blue-tilted tensor spectrum, and the other is for the almost scale-invariant spectrum. Using those configurations, we have computed the self-correlation function of the SGWB anisotropies and the cross-correlation function of those with the CMB temperature fluctuations. Then, we evaluated the possible maximum values of the amplitudes of those correlation functions in light of the backreaction constraint and the perturbativity condition. As a result, we have found that both anisotropies can dominate those originating from the late-time metric perturbations, so-called SW effects, and shown that the anisotropies with the maximum amplitudes can potentially be tested with PTA experiments. We have also found that the SNR for the scale-invariant case is much smaller than unity. However, since how large the SNR can be depends on the magnitude of $\bar{\Omega}_{\rm GW}$, those anisotropies could be tested with DECIGO/BBO as well if a relatively large tensor power spectrum compared to the scale-invariant case is detected at the deci-Hz frequency band. Here, our analysis of the detectability is based on order estimations, and hence it would be important to compute the noise spectrum explicitly and discuss the detectability in detail. We will leave it to our future work.
Acknowledgements
We thank Chunshan Lin and Giorgio Orlando for fruitful discussions. This work is supported by the grant No. UMO-2021/42/E/ST9/00260 from the National Science Centre, Poland (S.A.), MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe”, No. JP21H05189 (S.A), and JSPS KAKENHI Grants No. JP20K03968 (S.Y.), JP23H00108 (S.Y.) and JP24K00627 (S.Y.).
Appendix A Perturbed actions
A.1 Quadratic actions
By substituting the perturbed metric Eq. (2.2) into the action Eq. (2.1) and expanding up to the quadratic order, the quadratic action has been obtained as [82]
$S^{(2)}$ $=\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}-3\mathcal{G}_{T}\dot{\zeta}^{2}+\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{i}\zeta)^{2}+\Sigma\alpha^{2}-2\Theta\alpha\frac{\partial^{2}\beta}{a^{2}}+2\mathcal{G}_{T}\dot{\zeta}\frac{\partial^{2}\beta}{a^{2}}+6\Theta\alpha\dot{\zeta}-2\mathcal{G}_{T}\alpha\frac{\partial^{2}\zeta}{a^{2}}$
$\quad\quad\quad\quad\quad\quad\ +\frac{\mathcal{G}_{T}}{8}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr{]},$ (A.1)
$\Theta$ $:=-\dot{\phi}XG_{3X}+2HG_{4}-8HXG_{4X}-8HX^{2}G_{4XX}+\dot{\phi}G_{4\phi}+2X\dot{\phi}G_{4\phi{X}}$
$\quad-H^{2}\dot{\phi}(5XG_{5X}+2X^{2}G_{5XX})+2HX(3G_{5\phi}+2XG_{5\phi{X}}),$ (A.2)
$\Sigma$ $:=XG_{2X}+2X^{2}G_{2XX}+12H\dot{\phi}XG_{3X}+6H\dot{\phi}X^{2}G_{3XX}-2XG_{3\phi}-2X^{2}G_{3\phi{X}}-6H^{2}G_{4}$
$\quad+6\bigl{[}H^{2}(7XG_{4X}+16X^{2}G_{4XX}+4X^{3}G_{4XXX})-H\dot{\phi}(G_{4\phi}+5XG_{4\phi{X}}+2X^{2}G_{4\phi{X}X})\bigr{]}$
$\quad+2H^{3}\dot{\phi}\left(15XG_{5X}+13X^{2}G_{5XX}+2X^{3}G_{5XXX}\right)-6H^{2}X(6G_{5\phi}+9XG_{5\phi{X}}+2X^{2}G_{5\phi{X}X}).$ (A.3)
By varying the quadratic action with respect to $\alpha$ and $\beta$, one obtains the solutions of the constraint equations as
$\alpha$ $=\frac{\mathcal{G}_{T}}{\Theta}\dot{\zeta},$ (A.4)
$\beta$ $=\frac{1}{a\mathcal{G}_{T}}\biggl{(}a^{3}\mathcal{G}_{S}\psi-\frac{a\mathcal{G}_{T}^{2}}{\Theta}\zeta\biggr{)},$ (A.5)
where $\psi:=\partial^{-2}\dot{\zeta}$. Eliminating $\alpha$ and $\beta$ in the quadratic actions by using the solutions of the constraint equations, one obtains the quadratic actions in Eq. (2.5) and Eq. (2.6) where the coefficients are of the form,
$\mathcal{G}_{T}$ $:=2\left[G_{4}-2XG_{4X}-X(H\dot{\phi}G_{5X}-G_{5\phi})\right],$ (A.6)
$\mathcal{F}_{T}$ $:=2\left[G_{4}-X(\ddot{\phi}G_{5X}+G_{5\phi})\right],$ (A.7)
$\mathcal{G}_{S}$ $:=\mathcal{G}_{T}\biggl{(}\frac{\Sigma\mathcal{G}_{T}}{\Theta^{2}}+3\biggr{)},$ (A.8)
$\mathcal{F}_{S}$ $:=\frac{1}{a}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{a\mathcal{G}_{T}^{2}}{\Theta}\biggr{)}-\mathcal{F}_{T}.$ (A.9)
A.2 Cubic action from scalar-tensor-tensor interactions
Similarly to the quadratic actions, after expanding the Horndeski action with respect to metric perturbations up to the cubic order, the scalar-tensor-tensor type cubic action can be obtained as [101]
$S_{shh}$ $=\int{\rm d}t{\rm d}^{3}xa^{3}\biggl{[}\frac{3\mathcal{G}_{T}}{8}\zeta\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{8a^{2}}\zeta(\partial_{k}h_{ij})^{2}-\frac{\mu}{4}\dot{\zeta}\dot{h}_{ij}^{2}-\frac{\Gamma}{8}\alpha\dot{h}_{ij}^{2}-\frac{\mathcal{G}_{T}}{8a^{2}}\alpha(\partial_{k}h_{ij})^{2}$
$\quad\quad\quad\quad\quad\quad\ -\frac{\mu}{2a^{2}}\alpha\dot{h}_{ij}\partial^{2}h_{ij}-\frac{\mathcal{G}_{T}}{4a^{2}}\partial_{k}\beta\dot{h}_{ij}\partial_{k}h_{ij}-\frac{\mu}{2a^{2}}\biggl{(}\partial_{i}\partial_{j}\beta\dot{h}_{ik}\dot{h}_{jk}-\frac{1}{2}\partial^{2}\beta\dot{h}_{ij}^{2}\biggr{)}\biggr{]},$ (A.10)
where
$\Gamma$ $:=2G_{4}-8XG_{4X}-8X^{2}G_{4XX}-2H\dot{\phi}(5XG_{5X}+2X^{2}G_{5XX})+2X(3G_{5\phi}+2XG_{5\phi X}),$ (A.11)
$\quad-12HX(6G_{5\phi}+9XG_{5\phi X}+2X^{2}G_{5\phi XX}),$ (A.12)
$\mu$ $:=\dot{\phi}XG_{5X}.$ (A.13)
After eliminating $\alpha$ and $\beta$, one obtains
$S^{(3)}_{shh}=\int{\rm d}t{\rm d}^{3}x\biggl{[}a^{3}\mathcal{L}_{shh}+E_{shh}\biggr{]},$ (A.14)
where $\mathcal{L}_{shh}$ is given in Eq. (3.9), and $E_{shh}$ is defined by
$E_{shh}=\frac{\mu}{4\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}E^{s}+\frac{\mathcal{G}_{T}^{2}}{2\Theta\mathcal{F}_{T}}\biggl{(}\frac{\zeta}{2}+\frac{\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr{)}\dot{h}_{ij}E^{h}_{ij}.$ (A.15)
Each coefficient of the terms in $\mathcal{L}_{shh}$ takes the following form,
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl{\{}1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl{[}-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}\biggr{]}\biggr{\}},$ (A.16)
$b_{2}$ $=\frac{\mathcal{F}_{S}}{8},$ (A.17)
$b_{3}$ $=-\frac{\mathcal{G}_{S}}{4},$ (A.18)
$b_{4}$ $=\frac{\mathcal{G}_{T}}{8\Theta\mathcal{F}_{T}}(\mathcal{G}_{T}^{2}-\Gamma\mathcal{F}_{T})+\frac{\mu}{4}\biggl{[}\frac{\mathcal{G}_{S}}{\mathcal{G}_{T}}-1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}(6+g_{S})\biggr{]}+\frac{\mathcal{G}_{T}^{2}}{4}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mu}{\Theta\mathcal{F}_{T}}\biggr{)},$ (A.19)
$b_{5}$ $=\frac{\mu\mathcal{G}_{T}}{4\Theta}\biggl{(}\frac{\mathcal{F}_{S}\mathcal{G}_{T}}{\mathcal{F}_{T}\mathcal{G}_{S}}-1\biggr{)},$ (A.20)
$b_{6}$ $=-\frac{\mu\mathcal{G}_{S}}{2\mathcal{G}_{T}},$ (A.21)
$b_{7}$ $=\frac{\mu\mathcal{G}_{T}}{2\Theta}.$ (A.22)
The terms proportional to $E^{s}$ and $E^{h}_{ij}$ can be eliminated via the following field definition:
$h_{ij}$ $\to h_{ij}+\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\biggl{(}\zeta+\frac{2\mu}{\mathcal{G}_{T}}\dot{\zeta}\biggr{)}\dot{h}_{ij},$ (A.23)
$\zeta$ $\to\zeta+\frac{\mu}{8\mathcal{G}_{S}}\frac{\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}\dot{h}_{ij}^{2}.$ (A.24)
Appendix B Primordial STT bispectrum
In this section, we summarize the details of the computations of the primordial STT bispectrum. To obtain a general expression in the context of the non-Bunch-Davies states, we assume that both scalar and tensor modes are in non-Bunch-Davies states. In the main text, we chose $\eta_{0}$ in such a way that the primordial GWs at the small scales avoid the Trans-Planckian problem. However, the results obtained under Eq. (3.1) are valid at least for the primordial non-Gaussianities on CMB scales. Thus, to obtain a general expression that can be used widely, we consider possible values of $\eta_{0}$. We first assume that (i) the short modes of the tensor perturbations were initially in excited states, i.e., $\eta_{0}$ denotes the time when both long and short modes are on subhorizon scales, $-k_{i}\eta_{0}\gg 1$ (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{i}^{\prime}\eta_{0}\gg 1$ under the squeezed limit). Then, we consider the case in which (ii) the perturbations got excited from the Bunch-Davies state when the long mode crossed the sound horizon, $-c_{s}k_{1}\eta_{0}=1$ (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}=1$ under the squeezed limit). We also consider the case where (iii) the long mode was already on superhorizon scales, $-k_{L}\eta_{0}\ll 1$, while the short modes were still on subhorizon scales, $-k_{S}\eta_{0}\gg 1$, when the perturbations got excited (i.e., $-K^{\prime}\eta_{0}\gg 1,-K_{2}^{\prime}\eta_{0}\simeq-K_{3}^{\prime}\eta_{0}\ll 1$ under the squeezed limit). Last, we assume that (iv) both long and short modes were on superhorizon scales, $-k_{i}\eta_{0}\ll 1$, when the perturbations got excited (i.e., $-K^{\prime}\eta_{0}\ll 1,-K_{i}^{\prime}\eta_{0}\ll 1$ under the squeezed limit).
Let us parametrize the STT bispectrum as
$\mathcal{B}_{shh}^{s_{2}s_{3}}$ $={\rm Re}\Biggl{[}\sum^{7}_{i=1}(\alpha_{k_{1}}-\beta_{k_{1}})\left(\alpha^{(s_{2})}_{k_{2}}-\beta^{(s_{2})}_{k_{2}}\right)\left(\alpha^{(s_{3})}_{k_{3}}-\beta^{(s_{3})}_{k_{3}}\right)\frac{2}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\frac{H^{6}}{\mathcal{F}_{S}\mathcal{F}_{T}^{2}c_{s}c_{h}^{2}}b_{i}\mathcal{F}_{b_{i}}\mathcal{V}_{b_{i}}$
$\quad\quad\quad\quad\ +(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\Biggr{]}.$ (B.1)
B.1 $b_{1}$ term
By using the in-in formalism, the results for the $b_{1}$ term can be written as
$\mathcal{F}_{b_{1}}$ $=\frac{c_{h}^{4}k_{2}^{2}k_{3}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{1},K^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{1}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad+\mathcal{I}^{*}_{b_{1},K_{1}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{2}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{2}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.2)
and
$\mathcal{V}_{b_{1}}=\frac{1}{16k_{2}^{2}k_{3}^{2}}[k_{1}^{2}-(s_{2}k_{2}+s_{3}k_{3})^{2}]^{2},$ (B.3)
where
$\mathcal{I}_{b_{1},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (B.4)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (B.5)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (B.6)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.7)
Here, only $\mathcal{I}_{b_{1},K^{\prime}}$ appears in the case of the Bunch-Davies state, and thus one can evaluate the corrections to the bispectrum due to the excited modes by computing $\mathcal{I}_{b_{1},K^{\prime}}/\mathcal{I}_{b_{1},K_{i}^{\prime}}$.
In the case of (i), we obtain
$\mathcal{I}_{b_{1},K^{\prime}}$ $=\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}),$ (B.8)
$\mathcal{I}_{b_{1},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(-k_{1},k_{2},k_{3}),$ (B.9)
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (B.10)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (B.11)
where
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}}.$ (B.12)
In the cases of (ii), (iii), and (iv), $\mathcal{I}_{b_{1},K^{\prime}}$ and $\mathcal{I}_{b_{1},K_{1}^{\prime}}$ take the same form with Eq. (B.8) and Eq. (B.9), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{1},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{1}}(k_{1},-k_{2},k_{3}),$ (B.13)
$\mathcal{I}_{b_{1},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{1}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{1}}(k_{1},k_{2},-k_{3}),$ (B.14)
where
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}).$ (B.15)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
B.2 $b_{2}$ term
The results for the $b_{2}$ term can be written as
$\mathcal{F}_{b_{2}}$ $=\frac{1}{H^{2}}\biggl{[}\mathcal{I}_{b_{2},K^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{1}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\ +\mathcal{I}^{*}_{b_{2},K_{1}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{2}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{2}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\ +\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.16)
and
$\mathcal{V}_{b_{2}}=\frac{k_{1}^{2}-k_{2}^{2}-k_{3}^{2}}{2}\mathcal{V}_{b_{1}},$ (B.17)
where
$\mathcal{I}_{b_{2},K^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (B.18)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(-i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (B.19)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(-i+c_{h}k_{2}\eta)(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (B.20)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\frac{1}{\eta^{2}}(i+c_{s}k_{1}\eta)(i+c_{h}k_{2}\eta)(-i+c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta},$ (B.21)
In the case of (i), we obtain
$\mathcal{I}_{b_{2},K^{\prime}}$ $=\mathcal{J}_{b_{2}}(k_{1},k_{2},k_{3}),$ (B.22)
$\mathcal{I}_{b_{2},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(-k_{1},k_{2},k_{3}),$ (B.23)
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (B.24)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (B.25)
where
$\mathcal{J}_{b_{2}}(k_{1},k_{2},k_{3})$ $:=-\frac{1}{K^{\prime 2}}[c_{s}^{3}k_{1}^{3}+2c_{s}^{2}c_{h}k_{1}^{2}(k_{2}+k_{3})+2c_{s}c_{h}^{2}k_{1}(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})$
$\quad\quad\quad\quad\ +c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})],$ (B.26)
and we used the fact that the term proportional to $1/\eta_{*}$ with $\eta_{*}(\to 0)$ denoting the conformal time at the end of inflation does not contribute to the three-point function because of the following equation,
${\rm Re}\biggl{[}i(\alpha_{k_{1}}-\beta_{k_{1}})\left(\alpha^{(s_{2})}_{k_{2}}-\beta^{(s_{2})}_{k_{2}}\right)\left(\alpha^{(s_{3})}_{k_{3}}-\beta^{(s_{3})}_{k_{3}}\right)(\alpha^{*}_{k_{1}}-\beta^{*}_{k_{1}})\left(\alpha^{(s_{2})*}_{k_{2}}-\beta^{(s_{2})*}_{k_{2}}\right)$
×$\quad\ \times\left(\alpha^{(s_{3})*}_{k_{3}}-\beta^{(s_{3})*}_{k_{3}}\right)\biggr{]}=0.$ (B.27)
In the other three cases, $\mathcal{I}_{b_{2},K^{\prime}}$ and $\mathcal{I}_{b_{2},K_{1}^{\prime}}$ take the same form with Eq. (B.22) and Eq. (B.23), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{2},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{2}}(k_{1},-k_{2},k_{3}),$ (B.28)
$\mathcal{I}_{b_{2},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{2}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{2}}(k_{1},k_{2},-k_{3}),$ (B.29)
where
$\mathcal{S}_{b_{2}}(k_{1},k_{2},k_{3})$ $:=e^{iK^{\prime}\eta_{0}}\biggl{[}\frac{c_{h}}{K^{\prime 2}}(c_{h}^{2}k_{2}k_{3}(k_{2}+k_{3})+c_{s}^{2}k_{1}^{2}(k_{2}+k_{3})+c_{s}c_{h}k_{1}(k_{2}^{2}+4k_{2}k_{3}+k_{3}^{2}))$
$\quad\quad\quad\quad\ +\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr{]}.$ (B.30)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
B.3 $b_{3}$ term
The results for the $b_{3}$ term can be written as
$\mathcal{F}_{b_{3}}$ $=\frac{c_{h}^{2}c_{s}^{2}k_{1}^{2}k_{2}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{3},K^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{1}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\ +\mathcal{I}^{*}_{b_{3},K_{1}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{2}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{2}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.31)
and
$\mathcal{V}_{b_{3}}=-\frac{k_{1}^{2}-k_{2}^{2}+k_{3}^{2}}{2k_{1}^{2}}\mathcal{V}_{b_{1}},$ (B.32)
where
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK^{\prime}\eta},$ (B.33)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{1}^{\prime}\eta},$ (B.34)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\int{\rm d}\eta(i+c_{h}k_{3}\eta)e^{iK_{2}^{\prime}\eta},$ (B.35)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\int{\rm d}\eta(i-c_{h}k_{3}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.36)
In the case of (i), we obtain
$\mathcal{I}_{b_{3},K^{\prime}}$ $:=\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}),$ (B.37)
$\mathcal{I}_{b_{3},K_{1}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(-k_{1},k_{2},k_{3}),$ (B.38)
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (B.39)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (B.40)
where
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}}.$ (B.41)
In the other three cases, $\mathcal{I}_{b_{3},K^{\prime}}$ and $\mathcal{I}_{b_{3},K_{1}^{\prime}}$ take the same form with Eq. (B.37) and Eq. (B.38), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{3},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{3}}(k_{1},-k_{2},k_{3}),$ (B.42)
$\mathcal{I}_{b_{3},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{3}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{3}}(k_{1},k_{2},-k_{3}),$ (B.43)
where
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3}):=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}).$ (B.44)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
B.4 $b_{4}$ term
The results for the $b_{4}$ term can be written as
$\mathcal{F}_{b_{4}}$ $=\frac{c_{h}^{4}c_{s}^{2}k_{1}^{2}k_{2}^{2}k_{3}^{2}}{H}\biggl{[}\mathcal{I}_{b_{4},K^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{1}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\quad\ +\mathcal{I}^{*}_{b_{4},K_{1}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{2}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{2}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.45)
and
$\mathcal{V}_{b_{4}}=\mathcal{V}_{b_{1}},$ (B.46)
where
$\mathcal{I}_{b_{4},K^{\prime}}$ $:=-i\int{\rm d}\eta\eta^{2}e^{iK^{\prime}\eta},$ (B.47)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{1}^{\prime}\eta},$ (B.48)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{2}^{\prime}\eta},$ (B.49)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=i\int{\rm d}\eta\eta^{2}e^{iK_{3}^{\prime}\eta}.$ (B.50)
In the case of (i), we have
$\mathcal{I}_{b_{4},K^{\prime}}$ $=\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}),$ (B.51)
$\mathcal{I}_{b_{4},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(-k_{1},k_{2},k_{3}),$ (B.52)
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (B.53)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (B.54)
where
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3}):=\frac{2}{K^{\prime 3}}.$ (B.55)
In the other three cases, $\mathcal{I}_{b_{4},K^{\prime}}$ and $\mathcal{I}_{b_{4},K_{1}^{\prime}}$ take the same form with Eq. (B.51) and Eq. (B.52), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{4},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{4}}(k_{1},-k_{2},k_{3}),$ (B.56)
$\mathcal{I}_{b_{4},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{4}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{4}}(k_{1},k_{2},-k_{3}),$ (B.57)
where
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3}):=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2})$ (B.58)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
B.5 $b_{5}$ term
The results for the $b_{5}$ term can be written as
$\mathcal{F}_{b_{5}}$ $=c_{h}^{4}k_{1}^{2}k_{2}^{2}k_{3}^{2}\biggl{[}\mathcal{I}_{b_{5},K^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{1}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\ +\mathcal{I}^{*}_{b_{5},K_{1}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{2}^{\prime}}\alpha^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{2}^{\prime}}\beta^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}$
$\quad\quad\quad\quad\quad\ +\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{*}_{k_{1}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{3}^{\prime}}\beta^{*}_{k_{1}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}\biggr{]},$ (B.59)
and
$\mathcal{V}_{b_{5}}=k_{1}^{2}\mathcal{V}_{b_{1}},$ (B.60)
where
$\mathcal{I}_{b_{5},K^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK^{\prime}\eta},$ (B.61)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $:=-\int{\rm d}\eta\eta^{2}(-i+c_{s}k_{1}\eta)e^{iK_{1}^{\prime}\eta},$ (B.62)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{2}^{\prime}\eta},$ (B.63)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=\int{\rm d}\eta\eta^{2}(i+c_{s}k_{1}\eta)e^{iK_{3}^{\prime}\eta}.$ (B.64)
In the case of (i), we have
$\mathcal{I}_{b_{5},K^{\prime}}$ $=\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}),$ (B.65)
$\mathcal{I}_{b_{5},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(-k_{1},k_{2},k_{3}),$ (B.66)
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (B.67)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (B.68)
where
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3}):=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}.$ (B.69)
In the other three cases, $\mathcal{I}_{b_{5},K^{\prime}}$ and $\mathcal{I}_{b_{5},K_{1}^{\prime}}$ take the same form with Eq. (B.65) and Eq. (B.66), respectively, and we can write the results of the other two integrals as
$\mathcal{I}_{b_{5},K_{2}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{5}}(k_{1},-k_{2},k_{3}),$ (B.70)
$\mathcal{I}_{b_{5},K_{3}^{\prime}}$ $:=-\mathcal{J}_{b_{5}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{5}}(k_{1},k_{2},-k_{3}),$ (B.71)
where
$\mathcal{S}_{b_{5}}(k_{1},k_{2},k_{3}):=e^{iK^{\prime}\eta_{0}}\biggl{\{}\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}}+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr{\}}.$ (B.72)
In the case of (ii), $\eta_{0}$ is replaced with $-1/(c_{s}k_{1})$.
B.6 $b_{6}$ term
The results for the $b_{6}$ term can be written as
$\mathcal{F}_{b_{6}}=\mathcal{F}_{b_{4}},$ (B.73)
and
$\mathcal{V}_{b_{6}}=\frac{K}{32k_{1}^{2}k_{2}^{2}k_{3}^{2}}(k_{1}-k_{2}-k_{3})(k_{1}+k_{2}-k_{3})(k_{1}-k_{2}+k_{3})[k_{1}^{2}-(s_{2}k_{2}+s_{3}k_{3})^{2}],$ (B.74)
where $K=k_{1}+k_{2}+k_{3}$.
B.7 $b_{7}$ term
The results for the $b_{7}$ term can be written as
$\mathcal{F}_{b_{7}}=\mathcal{F}_{b_{5}},$ (B.75)
and
$\mathcal{V}_{b_{7}}=k_{1}^{2}\mathcal{V}_{b_{6}}.$ (B.76)
B.8 Field redefinition
Since the field redefinition does not bring any contributions from the far past to the resultant bispectrum, let us consider the superhorizon scales on which we have
$\dot{h}^{(s)}_{{\bf k}}\simeq ic_{h}^{2}k^{2}\eta^{2}Hh^{(s)}_{\bf k}.$ (B.77)
The quadratic-order terms in the field redefinition involve at least one time-derivative of the tensor perturbations, which means that those do not contribute to the bispectrum evaluated at the end of inflation $\eta\to 0$ (i.e., $\dot{h}^{(s)}_{{\bf k}}\to 0$). Therefore, the field redefinition always yields subleading contributions.
Appendix C $\mathcal{V}_{b_{i}}$ at the squeezed limit
One can take the squeezed limit to $\mathcal{V}_{b_{i}}$ straightforwardly as
$\mathcal{V}_{b_{1}}$ $=\frac{1}{16}\left(s_{2}+s_{3}\right)^{4}-\frac{1}{8}\left(s_{2}+s_{3}\right)^{3}\left(s_{2}-s_{3}\right)(\hat{n}\cdot\hat{q})\frac{q}{k}$
$\quad\ -\frac{1}{8}\left(s_{2}+s_{3}\right)^{2}\left[1-(\hat{n}\cdot\hat{q})^{2}\left(s_{2}^{2}-s_{2}s_{3}+s_{3}^{2}\right)\right]\frac{q^{2}}{k^{2}}+O((q/k)^{3}),$ (C.1)
$\mathcal{V}_{b_{2}}$ $=-\frac{1}{16}\left(s_{2}+s_{3}\right)^{2}k^{2}$
×$\quad\times\biggl{\{}\left(s_{2}+s_{3}\right)^{2}-2(\hat{n}\cdot\hat{q})(s_{2}-s_{3})(s_{2}+s_{3})\frac{q}{k}-\biggl{[}8+(1-8(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+(1-8(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}$
$\quad\quad\ +2s_{2}(s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3})\frac{q^{2}}{4k^{2}}\biggr{]}+O((q/k)^{3})\biggr{\}},$ (C.2)
$\mathcal{V}_{b_{3}}$ $=-\frac{(\hat{n}\cdot\hat{q})}{16}\frac{k}{q}\left(s_{2}+s_{3}\right)^{2}$
×$\quad\times\left\{\left(s_{2}+s_{3}\right)^{2}+\Big{[}s_{2}-4(\hat{n}\cdot\hat{q})^{2}s_{2}+s_{3}+4(\hat{n}\cdot\hat{q})^{2}s_{3}\Big{]}(s_{2}+s_{3})\frac{q}{2k(\hat{n}\cdot\hat{q})}\right.$ (C.3)
$\quad\quad\ \left.-\Big{[}2+(1-2(\hat{n}\cdot\hat{q})^{2})s_{2}^{2}+2(\hat{n}\cdot\hat{q})^{2}s_{2}s_{3}-(1+2(\hat{n}\cdot\hat{q})^{2})s_{3}^{2}\Big{]}\frac{q^{2}}{k^{2}}+O((q/k)^{3})\right\},$ (C.4)
$\mathcal{V}_{b_{4}}$ $=\mathcal{V}_{b_{1}},$ (C.5)
$\mathcal{V}_{b_{5}}$ $=q^{2}\mathcal{V}_{b_{1}},$ (C.6)
$\mathcal{V}_{b_{6}}$ $=\frac{1}{8}[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}+s_{3})^{2}-(\hat{n}\cdot\hat{q})[1-(\hat{n}\cdot\hat{q})^{2}](s_{2}^{2}-s_{3}^{2})\frac{q}{8k}$ (C.7)
$\quad-[1-(\hat{n}\cdot\hat{q})^{2}]\Big{\{}4+[1-4(\hat{n}\cdot\hat{q})^{2}]s_{2}^{2}+2[1-(\hat{n}\cdot\hat{q})^{2}]s_{2}s_{3}+[1-4(\hat{n}\cdot\hat{q})^{2}]s_{3}^{2}\Big{\}}\frac{q^{2}}{32k^{2}}$
$\quad+O((q/k)^{3}),$ (C.8)
$\mathcal{V}_{b_{7}}$ $=q^{2}\mathcal{V}_{b_{6}}.$ (C.9)
We have used the notation in Sec. 4.1. Also, $q$ and $k$ denote the wavenumber of the long and short modes, respectively. Based on these expressions, one can find that in the case with $s_{2}=-s_{3}$, $\mathcal{V}_{b_{6}}$ is suppressed at most in the order of $q/k$ and the others are more suppressed. Therefore, $\mathcal{V}_{b_{i}}$ take their maximum values for $s_{2}=s_{3}$.